Does training in study skills make a difference for student success? 8 students received training (M1= 85, SS1= 100) and 8 students did not (M2= 80, SS2= 124). Test using α = .05  

s(M1-M2) =

8

2

4

1
Answer #1

Answer: 4 Explanation: S(M1-M2) = SS1+SS2 v ni+n2-2 100+124 = 8+8-2 V16 S(M1-M2) = 4